OpenStudy (anonymous):

prove the identity: (sec^2 theda)(csc^2theda) = (csc^2theda) + (sec^2theda) PLEASE HELP!

OpenStudy (anonymous):



OpenStudy (anonymous):

the picture is cut off

OpenStudy (anonymous):

the answer is sec^2 x + csc^2 x, can u see it?

OpenStudy (anonymous):

I can see part of the answer

OpenStudy (anonymous):

tan^2 x +1 = sec^2 x 1 +cot^2 x = csc^2 x so, (sec^2x)(csc^2 x) =(tan^2 x +1) (1+ cot^2 x) = tan^2 x + (tan^2 x)(cot^2 x) +1 + cot^2 x = tan^2 x + 1 + 1 + cot^2 x = sec^2 x + csc^2 [(tan^2 x)(cot^2 x) = 1]

OpenStudy (anonymous):

thank you! I understand your steps. can you maybe help me with another one?

OpenStudy (anonymous):

no problem